# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v2

@@
Abstract
Multi-objective optimization (MOO) in deep learning aims to simultaneously optimize multiple conflicting objectives, a challenge frequently encountered in areas like multi-task learning and multi-criteria learning. Recent advancements in gradient-based MOO methods have enabled the discovery of diverse types of solutions, ranging from a single balanced solution to finite or even infinite Pareto sets, tailored to user needs. These developments have broad applications across domains such as reinforcement learning, computer vision, recommendation systems, and large language models. This survey provides the first comprehensive review of gradient-based MOO in deep learning, covering algorithms, theories, and practical applications. By unifying various approaches and identifying critical challenges, it serves as a foundational resource for driving innovation in this evolving field. A comprehensive list of MOO algorithms in deep learning is available at https://github.com/Baijiong-Lin/Awesome-Multi-Objective-Deep-Learning.
Index Terms:
@@ Abstract
1 Introduction
Multi-Objective Optimization (MOO) [1], which aims to optimize multiple potentially conflicting objectives simultaneously, has become an important topic in deep learning. A multi-objective optimization problem can be formulated as:
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$ (1)
@@ 1 Introduction
where $m\geq 2$ is the number of objectives, ${\bm{\theta}}$ is the decision variable and ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ is the feasible set of the decision variable.
@@
Figure 1: An overview of multi-objective deep learning.
Different from traditional single-objective optimization, which focuses on improving a single metric, MOO aims to obtain different solutions that achieve different trade-offs among objectives. Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex. MOO is expected to achieve a solution that matches the user-defined preference. Hence, compared to single-objective optimization, MOO is more practical for complex real-world applications.
Originally arising from operations research, MOO has been extensively studied across various disciplines of science and engineering [2]. In recent years, there has been a surge of interest in applying MOO techniques to deep learning, leading to the reformulation or extension of many learning paradigms as MOO problems.
In deep learning, there are generally two primary types of MOO scenarios. One scenario is multi-task learning [3], where a model is trained to handle multiple tasks simultaneously. For example, in computer vision, we may desire a single network to perform various dense prediction tasks concurrently, such as segmentation, depth estimation, and surface normal prediction [4]. In reinforcement learning [5], we might aim for a model that performs well on several robotic manipulation tasks. In the training of Large Language Models (LLMs), we expect the model to excel at different tasks, such as mathematics, reasoning, language translation, and summarization. However, these tasks often compete for shared network resources, leading to conflicts that can degrade performance on individual tasks. By formulating the problem as an MOO, where the performance on each task is treated as a separate objective [6], we can systematically manage these conflicts and optimize all task objectives simultaneously.
Another scenario involves incorporating additional evaluation criteria, such as fairness, safety, harmfulness mitigation, and efficiency, into learning paradigms. Considering these objectives is essential for practical real-world applications but may conflict with the primary goal of maximizing performance. For example, in LLMs training [7], we aim for models that not only perform well on tasks like language understanding but also adhere to safety standards, avoid producing harmful content, and ensure fairness across different users. In hardware-aware Neural Architecture Search (NAS) [8], we aim to discover network architectures that achieve high accuracy while also being efficient in terms of computational resources, energy consumption, or parameter size. However, reducing model size or computational complexity often leads to decreased performance. By treating such problems as MOO, where each criterion is an objective, we provide a structured way to explore these trade-offs.
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [ Gradient-based Multi-objective Deep Learning , fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA [9]; UW [10]; IMTL-L [11]; Lin et al. [12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SMG [20]; CR-MOGM [21]; MoCo [22]; PSMGD [23]; MoDo [24]; SDMGrad [25]; SGSMGrad [26]; IMTL-G [11]; Nash-MTL [27]; FairGrad [28]; UPGrad [29]; Aligned-MTL [30]; GradNorm [31]; DB-MTL [12]; FAMO [32], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [33]; GradVac [34]; GradDrop [35], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Methods Based on Preference Vectors (§4.1), finite [PMTL [36]; EPO [37, 38]; WC-MGDA [39]; PMGDA [40]; FERERO [41]; GMOOAR [42]; UMOD [43], finite_work] ] [Methods without Using Preference Vectors (§4.2), finite [GradHV [44, 45, 46, 47]; MOO-SVGD [48]; F4M [49]; SoM [50]; MosT [51], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [52]; CPMTL [53]; PHN-HVI [54]; SUHNPF [55]; Hyper-Trans [56], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [57]; YOTO [58]; Raychaudhuri et al. [59], infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL [60]; LORPMAN [61]; Chen and Kwok [62]; Panacea [63]; Tang et al. [64]; PaLoRA [65], infinite_work] ] ] ]
@@
Figure 2: Taxonomy of existing MOO algorithms in deep learning.
1.1 Taxonomy
Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches:
  • To obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization. These methods focus on adjusting the weight of each objective to achieve a solution that is not overly biased toward any one objective.
  • To obtain a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously. This allows users to select from a set of solutions that best meet their specific needs.
  • To generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones. Algorithms in this category propose efficient structures to learn this subspace. During training, preference vectors are randomly sampled in each iteration. Once a preference is sampled, the optimization reduces to solving for either a single solution or a finite set, making it possible to use techniques from the first two approaches.
Figure 2 illustrates the taxonomy of current MOO algorithms in deep learning, serving as a guide to the methods discussed in this paper. By exploring these avenues, MOO provides a comprehensive framework for addressing complex optimization problems in deep learning, enabling the development of models that can effectively balance multiple competing objectives.
1.2 Comparison with Related Surveys
Some surveys have attempted to cover aspects of MOO. For example, Zhang and Yang [3] provide a comprehensive overview of traditional multi-task methods within the machine learning domain. Crawshaw [66] offers a summary of multi-task methods in deep learning up to the year 2020. Yu et al. [67] present a comprehensive survey that covers multi-task learning from traditional machine learning to deep learning. Eichfelder [68] discusses advancements in MOO from a general optimization perspective, without specific emphasis on deep learning. Wei et al. [69] focus on evolutionary MOO algorithms, but these algorithms are usually not suitable for deep neural networks due to their large parameter spaces and computational demands. Peitz and Hotegni [70] survey the MOO algorithms for deep learning but cover only a limited selection of gradient-based MOO methods and do not include theory and applications. Other surveys concentrate on applications of MOO in areas such as dense prediction tasks [4], recommend systems [71], and natural language processing [72].
Different with the previous surveys, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications. To the best of our knowledge, it is the first survey paper to focus on the gradient-based MOO methods in deep learning.
TABLE I: Summary of Notations.
@@ 1.2 Comparison with Related Surveys
$\Delta_{m-1}$ $(m-1)$-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
[…]
@@
1.3 Organization of the Survey
The notations used in this survey are summarized in Table I. An overview of MOO in deep learning is provided in Figure 1. The rest of this survey is organized as follows:
  • Section 2 presents preliminaries in MOO, including key definitions and concepts;
  • Section 3 discusses methods for finding a single Pareto optimal solution, covering loss balancing and gradient balancing approaches;
  • Sections 4 and 5 focus on methods for identifying a finite and an infinite set of Pareto optimal solutions, respectively;
  • Section 6 delves into the theoretical analysis of convergence and generalization in gradient-based multi-objective optimization algorithms;
  • Section 7 showcases various applications in deep learning, including reinforcement learning, Bayesian optimization, computer vision, and large language models;
  • Section 8 offers useful resources for multi-objective deep learning, such as datasets and software tools;
  • Section 9 highlights challenges in the field and suggests promising directions for future research;
  • Section 10 summarizes this survey.
2 Preliminary: Key Concepts in MOO
In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to find Pareto solutions of MOO problems in defined in Section 2.2. An illustration of some Pareto concepts is in Appendix A.
2.1 Pareto Concepts
@@
Definition 1 ((strict) Pareto dominance [1])
@@
Definition 2 ((weak) Pareto optimality [1])
@@
Definition 3 (Pareto set (PS) and Pareto front (PF) [1])
@@
Definition 4 (Preference vector)
@@
Definition 5 (Pareto stationary [73])
@@
Pareto stationarity is a necessary condition for achieving Pareto optimality. If all objectives are convex with $\lambda_{i}>0,\forall i$, it also serves as the Karush-Kuhn-Tucker (KKT) sufficient condition for Pareto optimality [73].
When evaluating the performance of a set of obtained solutions, the Hypervolume (HV) [74] indicator is one of the most widely used performance indicators, which is defined as follows.
Definition 6 (Hypervolume [74])
@@
$\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\Lambda({{\bm{p}}\;|\;\exists{\bm{q}}\in{\mathbb{S}}:{\bm{q}}\preceq{\bm{p}}\preceq{\bm{r}}}),$ (2)
@@
HV provides a unary measure of both convergence and diversity of a solution set. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
2.2 Scalarization Methods
Linear scalarization. The most straightforward way to solve MOO problems is to reformulate them as single-objective optimization problems weighted by the given preference vector ${\bm{\alpha}}$ as follows:
$\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{LS}}({\bm{\theta}}):=\sum_{i=1}^{m}\alpha_{i}f_{i}({\bm{\theta}}).$ (3)
Linear scalarization is widely used in practice due to its simplicity. However, this method suffers from a limitation: it can only identify solutions on the convex part of the PF. For PFs with a concave shape, using LS only finds the endpoints of a PF. Tchebycheff scalarization. Another way to convert an MOP to a single one is to use the Tchebycheff scalarization function [75, 76], defined as:
$\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{Tche}}({\bm{\theta}}):=\max_{i\in[m]}\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*}),$ (4)

### v3

@@
Abstract.
Many modern deep learning applications require balancing multiple objectives that are often conflicting. Examples include multi-task learning, fairness-aware learning, and the alignment of Large Language Models (LLMs). This leads to multi-objective deep learning, which tries to find optimal trade-offs or Pareto-optimal solutions by adapting mathematical principles from the field of Multi-Objective Optimization (MOO). However, directly applying gradient-based MOO techniques to deep neural networks presents unique challenges, including high computational costs, optimization instability, and the difficulty of effectively incorporating user preferences. This paper provides a comprehensive survey of gradient-based techniques for multi-objective deep learning. We systematically categorize existing algorithms based on their outputs: (i) methods that find a single, well-balanced solution, (ii) methods that generate a finite set of diverse Pareto-optimal solutions, and (iii) methods that learn a continuous Pareto set of solutions. In addition to this taxonomy, the survey covers theoretical analyses, key applications, practical resources, and highlights open challenges and promising directions for future research. A comprehensive list of multi-objective deep learning algorithms is available at https://github.com/Baijiong-Lin/Awesome-Multi-Objective-Deep-Learning.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Machine learningccs: Computing methodologies Multi-task learning
1. Introduction
Traditional deep learning often focuses on optimizing a single learning objective, such as minimizing the prediction error or maximizing the likelihood. However, many real-world applications require balancing multiple, often conflicting, objectives. For instance, a computer vision system might need to perform the tasks of segmentation, depth estimation, and surface normal prediction simultaneously (Vandenhende et al., 2021), thus moving from single-task learning to multi-task learning (Zhang and Yang, 2022). Similarly, Large Language Models (LLMs) are expected to excel at diverse tasks like reasoning and coding, while also being safe, fair, and harmless (Wang et al., 2023b). Multi-Objective Optimization (MOO) (Miettinen, 1999), a field originating from operations research (Ehrgott, 2005), provides a formal framework for navigating these trade-offs. It has been widely studied across science and engineering, with diverse applications such as in finance (Steuer and Na, 2003), engineering design (Marler and Arora, 2004), and transportation planning (Tzeng et al., 2005).
Recently, there has been a surge of interests in adapting MOO for deep learning, re-framing many popular learning problems as multi-objective optimization problems. Examples include multi-task learning where each task performance is considered as an objective (Sener and Koltun, 2018), fairness-aware learning where accuracy is one objective and fairness metrics constitute the other objectives (Martinez et al., 2020), LLM alignment where each alignment criterion (such as helpfulness, harmlessness, and honesty) represents a separate objective (Wang et al., 2023b), and federated learning where the performance on each client is treated as an individual objective (Hu et al., 2022). All these problems can be formally expressed as:
(1) $\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~~{\bm{f}}({\bm{\theta}})$ $:=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$
@@ 1. Introduction
where $m\geq 2$ is the number of objectives, $\bm{\theta}$} is the decision variable, ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ is the feasible set of the decision variable, and each $f_{i}:{\mathcal{K}}\to{\mathbb{R}}$ represents an individual objective function to be minimized.
Unlike single-objective optimization, which focuses on finding a single best solution, MOO recognizes that no single solution is optimal for all objectives simultaneously. Instead, MOO aims to identify a set of solutions that represent different trade-offs between objectives, collectively known as the Pareto set (Miettinen, 1999). The Pareto set consists of all solutions where improving any objective would necessarily worsen at least one other objective, representing the best possible trade-offs available. In real-world applications, users often have varying preferences for these trade-offs. For instance, in the development of LLMs, a customer service application might emphasize harmlessness and safety, whereas an educational tool might prioritize reasoning and factual accuracy. To address this variability, users can specify their preferences using a vector, ${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}\in\Delta_{m-1}$, where $\Delta_{m-1}=\{{\bm{\alpha}}\in{\mathbb{R}}^{m}_{+}:\sum_{i=1}^{m}\alpha_{i}=1\}$ is the probability simplex and each $\alpha_{i}$ represents the importance assigned to the $i$-th objective. MOO approaches can accommodate these user-defined preferences, enabling the discovery of solutions that align with individual needs.
MOO methods can be broadly divided into gradient-free and gradient-based approaches. Gradient-free methods, such as evolutionary algorithms (Deb et al., 2002; Zhang and Li, 2007) and particle swarm optimization (Coello et al., 2004), explore the solution space using population-based sampling. While effective for traditional low-dimensional MOO problems (Deb, 2011), they struggle with deep neural networks, which often have millions or even billions of parameters. Due to the curse of dimensionality, the parameter space becomes too large to explore efficiently without gradient information, leading to slow or failed convergence. This makes gradient-based methods (Désidéri, 2012; Mukai, 1980; Fliege and Svaiter, 2000; Liu et al., 2021b) preferable for deep neural networks, as they efficiently guide the search using objective gradients.
However, gradient-based MOO faces several significant challenges in deep learning. Firstly, incorporating user preferences is difficult because directly weighting objectives often produces unaligned solutions. Secondly, training even one neural network is computationally expensive, posing a major challenge to efficiently approximating the entire Pareto set of networks. Finally, the use of mini-batch optimization in deep learning introduces high variance, which can further destabilize the optimization process. These challenges have motivated extensive research efforts to develop specialized MOO techniques for deep learning. This paper provides a comprehensive survey of these advances, systematically categorizing existing approaches and identifying key open problems in the field.
1.1. Comparison with Related Surveys on MOO
Some surveys have attempted to cover aspects of MOO. For example, Zhang and Yang (Zhang and Yang, 2022) provide a comprehensive overview of traditional multi-task methods within the machine learning domain. Crawshaw (Crawshaw, 2020) offers a summary of multi-task methods in deep learning up to the year 2020. Yu et al. (Yu et al., 2024) present a comprehensive survey that covers multi-task learning from traditional machine learning to deep learning. Eichfelder (Eichfelder, 2021) discusses advancements in MOO from a general optimization perspective, without specific emphasis on deep learning. Wei et al. (Wei et al., 2021) focus on evolutionary MOO algorithms, but these algorithms are usually not suitable for deep neural networks due to their large parameter spaces and computational demands. Peitz and Hotegni (Peitz and Hotegni, 2024) survey the MOO algorithms for deep learning but cover only a limited selection of gradient-based MOO methods and do not include theory and applications. Other surveys concentrate on applications of MOO in areas such as dense prediction tasks (Vandenhende et al., 2021), recommend systems (Zhang et al., 2023b), and natural language processing (Chen et al., 2024b).
Different with the previous surveys, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications, as the overview shown in Figure 1. To the best of our knowledge, it is the first survey paper focusing on the gradient-based MOO methods in deep learning.
@@ 1.1. Comparison with Related Surveys on MOO
Figure 1. An overview of gradient-based multi-objective deep learning.
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [ Gradient-based Multi-objective Deep Learning , fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA (Liu et al., 2019a); UW (Kendall et al., 2018); IMTL-L (Liu et al., 2021a); Lin et al. (Lin et al., 2023); MOML (Ye et al., 2021, 2024b); Auto-$\lambda$ (Liu et al., 2022); FORUM (Ye et al., 2024a); RW (Lin et al., 2022c); STCH (Lin et al., 2024b), single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA (Sener and Koltun, 2018); CAGrad (Liu et al., 2021b); SMG (Liu and Vicente, 2021); CR-MOGM (Zhou et al., 2022); MoCo (Fernando et al., 2023); PSMGD (Xu et al., 2025b); MoDo (Chen et al., 2023a); SDMGrad (Xiao et al., 2023); SGSMGrad (Zhang et al., 2025); IMTL-G (Liu et al., 2021a); Nash-MTL (Navon et al., 2022); FairGrad (Ban and Ji, 2024); UPGrad (Quinton and Rey, 2024); Aligned-MTL (Senushkin et al., 2023); GradNorm (Chen et al., 2018); DB-MTL (Lin et al., 2023); FAMO (Liu et al., 2023), single_work2] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad (Yu et al., 2020a); GradVac (Wang et al., 2021b); GradDrop (Chen et al., 2020), single_work2] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Methods Based on Preference Vectors (§4.1), finite [PMTL (Lin et al., 2019b); EPO (Mahapatra and Rajan, 2020, 2021); WC-MGDA (Momma et al., 2022); PMGDA (Zhang et al., 2024b); FERERO (Chen et al., 2024a); GMOOAR (Chen and Kwok, 2022); UMOD (Zhang et al., 2024a), finite_work] ] [Methods without Using Preference Vectors (§4.2), finite [GradHV (Deist et al., 2021, 2020; Emmerich et al., 2007; Wang et al., 2017); MOO-SVGD (Liu et al., 2021d); F4M (Lin et al., 2025c); SoM (Ding et al., 2024); MosT (Li et al., 2025), finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN (Navon et al., 2021); CPMTL (Lin et al., 2020); PHN-HVI (Hoang et al., 2023); SUHNPF (Gupta et al., 2022); Hyper-Trans (Tuan et al., 2024), infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS (Ruchte and Grabocka, 2021); YOTO (Dosovitskiy and Djolonga, 2020); Raychaudhuri et al. (Raychaudhuri et al., 2022), infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL (Dimitriadis et al., 2023); LORPMAN (Chen and Kwok, 2024a); Chen and Kwok (Chen and Kwok, 2024b); Panacea (Zhong et al., 2024); Tang et al. (Tang et al., 2024); PaLoRA (Dimitriadis et al., 2025), infinite_work] ] ] ]
@@
Figure 2. Taxonomy of existing gradient-based multi-objective deep learning algorithms.
1.2. Organization of the Survey
In this survey, we categorize gradient-based multi-objective deep learning algorithms based on the type of output they obtain: (Figure 2): (i) Algorithms that obtain a single, well-balanced model that is not overly biased toward any one objective: This is achieved through techniques like loss balancing and gradient balancing, which adjust the weight of each objective to resolve conflicts during optimization. (ii) Algorithms that obtain a finite set of Pareto-optimal solutions that allows users to select from multiple options based on their specific needs: This is accomplished by either dividing the problem into subproblems using preference vectors or directly optimizing for multiple solutions simultaneously. (iii) Algorithms that obtain a continuous set of Pareto-optimal models that can be generated on-demand for any given preference: This is achieved by learning a solution subspace using efficient structures, where preference vectors are randomly sampled during training and the optimization adapts techniques from single or finite solution methods.
The rest of this survey is organized as follows: Section 2 presents MOO preliminaries, including MOO definitions and concepts; Section 3 discusses methods for finding a single Pareto-optimal solution, covering loss balancing and gradient balancing approaches; Section 4 focuses on methods for identifying a set of finite Pareto-optimal solutions; Section 5 covers methods for identifying a set of infinite Pareto optimal solutions; Section 6 delves into the theoretical analysis of convergence and generalization in gradient-based multi-objective optimization algorithms; Section 7 showcases various applications in deep learning, including reinforcement learning, Bayesian optimization, computer vision, neural architecture search, recommender systems, and large language models; Section 8 offers useful resources on datasets and software tools for multi-objective deep learning; Section 9 highlights challenges in the field and suggests promising directions for future research; Section 10 summarizes this survey.
1.3. Notations
The notations used in this survey are summarized in Table 1. Scalars are represented by non-bold letters; vectors and matrices are denoted by boldface type. Subscripts are used as indices for elements, and superscripts typically denote iteration numbers or indices within a solution set.
Table 1. Summary of Notations.
@@ 1.3. Notations
$\Delta_{m-1}$ (m1)(m-1)-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
$\|\cdot\|$ $\ell_{2}$ norm.
@@
2. Preliminary: Multi-Objective Optimization
In this section, we first introduce some MOO concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to find Pareto solutions of MOO problems in defined in Section 2.2.
2.1. Key Concepts in Multi-Objective Optimization
@@
Definition 2.1 ((strict) Pareto dominance (Miettinen, 1999)).
@@
Definition 2.2 ((weak) Pareto optimality (Miettinen, 1999)).
@@
Definition 2.3 (Pareto set (PS) and Pareto front (PF) (Miettinen, 1999)).
@@
Figure 3 illustrates the Pareto concepts on the two-objective problem. The blue circles represent Pareto optimal solutions (defined in Definition 2.2), which indicates that no solution can improve one objective without worsening the other. The blue curve connecting the blue circles denotes the Pareto front, as defined in Definition 2.3. The yellow circles indicate weak Pareto optimal solutions since they can be improved in one objective without negatively impacting the other. For example, comparing ${\bm{\theta}}_{2}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{2}$ is a weak Pareto optimal solution since it can be improved in the second objective without affecting the first. The red circles represent solutions that are not Pareto optimal because there exists another solution that strictly Pareto dominates them. For instance, comparing ${\bm{\theta}}_{3}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{3}$ is not a Pareto optimal solution as ${\bm{\theta}}_{1}$ outperforms it in both objectives.
Figure 3. Illustration of Pareto optimality concepts. The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively. The blue curve represents the Pareto front.
Definition 2.4 (Preference vector).
@@
Definition 2.5 (Pareto stationary (Désidéri, 2012)).
@@
Pareto stationarity is a necessary condition for achieving Pareto optimality. If all objectives are convex with $\lambda_{i}>0,\forall i$, it also serves as the Karush-Kuhn-Tucker (KKT) sufficient and necessary condition for Pareto optimality (Désidéri, 2012). When evaluating the performance of a set of obtained solutions, the Hypervolume (HV) (Zitzler and Thiele, 1998) indicator is one of the most widely used performance indicators, which is defined as follows.
Definition 2.6 (Hypervolume (Zitzler and Thiele, 1998)).
@@
(2) $\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\Lambda({{\bm{p}}\;|\;\exists{\bm{q}}\in{\mathbb{S}}:{\bm{q}}\preceq{\bm{p}}\preceq{\bm{r}}}),$
@@
HV provides a unary measure of both convergence and diversity of a solution set. A larger HV indicates that the obtained solutions are more diverse and closer the PF in the objective space.
2.2. Scalarization Methods
Linear Scalarization (LS). The most straightforward way to solve MOO problems is to reformulate them as single-objective optimization problems weighted by the given preference vector $\bm{\alpha}$} as follows:
(3) $\min_{{\bm{\theta}}\in{\mathcal{K}}}g_{\bm{\alpha}}^{\mathrm{LS}}({\bm{\theta}}):=\sum_{i=1}^{m}\alpha_{i}f_{i}({\bm{\theta}}).$
LS is widely used in practice due to its simplicity. However, this method suffers from a limitation: it can only identify solutions on the convex part of the PF. For PFs with a concave shape, using LS only finds the endpoints of a PF.
Tchebycheff Scalarization. Another way to convert a MOO problem to a single objective optimization problem is to use the Tchebycheff scalarization function (Bowman Jr, 1976; Steuer and Choo, 1983), defined as:
(4) $\min_{{\bm{\theta}}\in{\mathcal{K}}}g_{\bm{\alpha}}^{\mathrm{Tche}}({\bm{\theta}}):=\max_{i\in[m]}\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*}),$